# FiRST: Finetuning Router-Selective Transformers for Input-Adaptive Latency Reduction

### v3

@@ 3 Proposed Solution: FiRST
Let us first understand what it entails to learn an optimal subsequence of layers for any input. Consider the full transformer sequence to be =$\mathcal{F}^{*}=\{\phi_{1},\phi_{2},\ldots,\phi_{m}\}$. Any optimal subsequence for an input $\mathcal{X}$: $\mathcal{F}_{OPT,\mathcal{X}}$ could be thought of as finding an optimal path through a binary tree of functions. Formally, let every level in the binary tree correspond to a transformer layer and the $0^{th}$ layer corresponds to the initial embedding look up; i.e., at depth $i\in[m]$, there would be $2^{i}$ nodes, each corresponding to either $\phi_{i}$ or $\overline{\phi_{i}}$, where the former denotes that a particular transformer layer is included in the optimal path whereas the latter denotes that it is not included. Each (of the $2^{i-1}$ nodes) $\phi_{i}$ or $\overline{\phi_{i}}$ has two children, corresponding to the next transformer layer: $\phi_{i+1}$ and $\overline{\phi_{i+1}}$ (See Figure 1). In such a tree structure, for example, the path $\{\phi_{i},\overline{\phi_{i+1}},\phi_{i+2}\}$ indicates the subsequence of transformer layers $\{\phi_{i},\phi_{i+2}\}$. For any transformer layer $\phi_{i}$ in this tree, let $Anc(\phi_{i})=k,0\leq k<i$ denote the the lowest ancestor node where the corresponding transformer node $\phi_{k}$ is included in the sequence. In the above example, $Anc(\phi_{i+2})=\phi_{i}$.
@@
$H^{j}_{i}$ =$=\begin{cases}H^{j}_{k},&\text{if }\phi_{i}\notin\mathcal{F},\\ H^{j}_{k}+\phi_{i}(H^{1}_{k},H^{2}_{k},\ldots,H^{j}_{k}),&\text{if }\phi_{i}\in\mathcal{F}.\end{cases}$ (3)
@@ 4.1 Adaptive Router Module
The adaptive router module is a single-layer neural network without bias, positioned before every layer in the model. During training of the router, all model parameters except the router weights remain frozen. For the first layer, it takes the tokenized input, and for each of the subsequent layers, it takes the output of the preceding layer as input. Mathematically speaking, for any layer $i$, given a batch of $B$ tokenized inputs sequences, where each sequence has $n$ tokens and is embedded in to $\mathbb{R}^{D}$, the adaptive router module takes as input a $B\times n\times D$ tensor output of layer $(i-1)$ and outputs a $B\times n\times 1$ tensor. Subsequently, corresponding to each value (or, token) in the $B\times n\times 1$ tensor, we apply a sigmoid function to ensure that all entries in the tensor are in the interval $[0,1]$. Following this, we take a mean operation at the sequence level - we take a mean of all the weights in a sequence to output a $B\times 1\times 1$ tensor. For each sequence in the batch, the corresponding entry is the probability $\rho_{i}$ with which the sequence passes through the layer $i$. The input sequence skips the layer $i$ with probability $1-\rho_{i}$. During training, the output of a layer is modified using a skip connection, incorporating the probability $\rho_{i}$ (see Figure: 2).
@@
  • Cross-entropy loss: Standard difference between actual and predicted probability distributions to ensure the quality of generation: $\mathcal{L}_{\mathrm{CE}}=-\sum_{x\in\mathcal{X}}\mathcal{Y}^{*}_{\mathcal{X}}\log(\hat{\mathcal{Y}})$.
  • Regularization loss: Adds a penalty term to reduce overfitting to noise: $\mathcal{L}_{\mathrm{Reg}}=\sum_{i\in[m]}||R_{i}||^{2}$, where $||R_{i}||^{2}$ denotes the $\ell_{2}$ norm of the router weights for the $i^{\text{th}}$ layer router, and there are $m$ layers in the model.
@@ 4.3 Inference for FiRST
Figure 2: Skip connection used for router training. With probability $p$, the sequence is processed by the layer and with probability $1-p$, the layer is skipped. During inference, routers make the decision of whether a sequence will skip a particular layer or pass through it.
@@ 5.2 Experimental Results on Different Tasks
CNN/DailyMail Dataset: For LLaMA-3-8B, at $\sim 15\%$ skipping, our method outperforms the Base + LoRA setting(Table: 3) while obtaining a 12% improvement in TPOT (Table 4). For LLaMA-3.2-3B, at $15\%$, the quality is comparable ($\sim 98\%$) to gold and other baselines with $12\%$ improvement in TPOT. At $24\%$, FiRST is significantly better than other layer skipping strategies, while achieving $>20\%$ improvement in latency.
@@
SQuAD Dataset: For the LLaMA-3-8B model, FiRST(at 10% skip level) maintains over 95% of the performance of the gold standard (Base + LoRA without skipping)(Table 2), with overall latency gains upto 16% (Table 4). It is significantly better in quality than all other baselines across all metrics for different levels of skipping. For LLaMA-3.2-3B, again FiRST is $>95\%$ in output quality of gold (base + LoRA) for 10% skipping (Table 2) with similar gains in latency (upto 16% overall) (Table 4) over the LoRA fine-tuned base model. Moreover, it is better than all other layer skipping strategies across all metrics.
Natural Questions Dataset: For the LLaMA-3-8B model, FiRST retains over $92\%$ of the EM achieved by the non-skipping gold standard (Table 2), with overall latency gains of 5-12% (Table 4). It is significantly better in quality than all other baselines for different levels of skipping. For LLaMA-3.2-3B, again FiRST is $>93\%$ in output quality of gold (base + LoRA) for 10% skipping (Table 2) with gains in latency of 4-10% overall (Table 4) over the LoRA fine-tuned base model. Moreover, it is better than all other layer skipping strategies. While more sophisticated retrieval strategies could further enhance performance, our goal was to demonstrate that FiRST maintains quality performance and latency gains even in a retrieval-based QA setting.
@@ A.4 Training and Inference Setup
[…]
@@ A.8 Generalizability and Reusability of Routers
We also investigate whether a router trained on one dataset can be applied directly to another dataset for the same task, without any additional training. Analysis of layer-skipping patterns suggests that models processing different datasets of the same task exhibit similar internal behavior, implying that a router learned on one dataset could be reused elsewhere. To validate this idea, we train routers on English→German translation data and test them on English→Chinese translation, and likewise train on SQuAD for question answering before evaluating on the Natural Questions dataset. The quantitative results of these cross-dataset evaluations are presented in Table 13. In the QA experiments, the router trained and tested on NQ itself (denoted as Original in Table 13) and the router trained on SQuAD but tested on NQ (Cross-dataset) both skip a comparable fraction of layers. Despite being trained on different data, the SQuAD-trained router achieves QA accuracy (EM score) very close to that of the NQ-trained router, as detailed in Table 13. This indicates that the essential decision patterns learned by the router transfer well across QA datasets when using a similar level of layer skipping. A similar pattern emerges in machine translation. Whether trained on English→German or directly on English→Chinese data, routers that skip the same proportion of layers produce very similar translation quality metrics (BLEU and COMET in Table 13) on the English→Chinese task. This holds across different model sizes and skip settings. Together, these results demonstrate that routers learned on one dataset can be effectively reused on another dataset for the same task, as long as they are configured to skip a similar amount of computation. This reusability can lead to substantial savings in both training time and computing resources.

### v4

@@ 3 Proposed Solution: FiRST
Let us first understand what it entails to learn an optimal subsequence of layers for any input. Consider the full transformer sequence to be $\mathcal{F}^{*}=\{\phi_{1},\phi_{2},\ldots,\phi_{m}\}$. Any optimal subsequence for an input $\mathcal{X}$: $\mathcal{F}_{OPT,\mathcal{X}}$ could be thought of as finding an optimal path through a binary tree of functions. Formally, let every level in the binary tree correspond to a transformer layer and the $0^{th}$ layer corresponds to the initial embedding look up; i.e., at depth $i\in[m]$, there would be $2^{i}$ nodes, each corresponding to either $\phi_{i}$ or $\overline{\phi_{i}}$, where the former denotes that a particular transformer layer is included in the optimal path whereas the latter denotes that it is not included. Each (of the $2^{i-1}$ nodes) $\phi_{i}$ or $\overline{\phi_{i}}$ has two children, corresponding to the next transformer layer: $\phi_{i+1}$ and $\overline{\phi_{i+1}}$ (See Figure 1). In such a tree structure, for example, the path $\{\phi_{i},\overline{\phi_{i+1}},\phi_{i+2}\}$ indicates the subsequence of transformer layers $\{\phi_{i},\phi_{i+2}\}$. For any transformer layer $\phi_{i}$ in this tree, let $Anc(\phi_{i})=k,0\leq k<i$ denote the the lowest ancestor node where the corresponding transformer node $\phi_{k}$ is included in the sequence. In the above example, $Anc(\phi_{i+2})=\phi_{i}$.
@@
$H^{j}_{i}$ $=\begin{cases}H^{j}_{k},&\text{if }\phi_{i}\notin\mathcal{F},\\ H^{j}_{k}+\phi_{i}(H^{1}_{k},H^{2}_{k},\ldots,H^{j}_{k}),&\text{if }\phi_{i}\in\mathcal{F}.\end{cases}$ (3)
@@ 4.1 Adaptive Router Module
The adaptive router module is a single-layer neural network without bias, positioned before every layer in the model. During training of the router, all model parameters except the router weights remain frozen. For the first layer, it takes the tokenized input, and for each of the subsequent layers, it takes the output of the preceding layer as input. Mathematically speaking, for any layer $i$, given a batch of $B$ tokenized inputs sequences, where each sequence has $n$ tokens and is embedded in to $\mathbb{R}^{D}$, the adaptive router module takes as input a $B\times n\times D$ tensor output of layer (i1)(i-1) and outputs a $B\times n\times 1$ tensor. Subsequently, corresponding to each value (or, token) in the $B\times n\times 1$ tensor, we apply a sigmoid function to ensure that all entries in the tensor are in the interval $[0,1]$. Following this, we take a mean operation at the sequence level - we take a mean of all the weights in a sequence to output a $B\times 1\times 1$ tensor. For each sequence in the batch, the corresponding entry is the probability $\rho_{i}$ with which the sequence passes through the layer $i$. The input sequence skips the layer $i$ with probability $1-\rho_{i}$. During training, the output of a layer is modified using a skip connection, incorporating the probability $\rho_{i}$ (see Figure: 2).
@@
[…]
@@ 4.3 Inference for FiRST
Figure 2: Skip connection used for router training. With probability $p$, the sequence is processed by the layer and with probability 1p1-p, the layer is skipped. During inference, routers make the decision of whether a sequence will skip a particular layer or pass through it.
@@ 5.2 Experimental Results on Different Tasks
CNN/DailyMail Dataset: For LLaMA-3-8B, at $\sim 15\%$ skipping, our method outperforms the Base + LoRA setting(Table: 3) while obtaining a 12% improvement in TPOT (Table 4). For LLaMA-3.2-3B, at $15\%$, the quality is comparable ($\sim 98\%$) to gold and other baselines with $12\%$ improvement in TPOT. At $24\%$, FiRST is significantly better than other layer skipping strategies, while achieving $20\%$ improvement in latency.
@@
SQuAD Dataset: For the LLaMA-3-8B model, FiRST(at 10% skip level) maintains over 95% of the performance of the gold standard (Base + LoRA without skipping)(Table 2), with overall latency gains upto 16% (Table 4). It is significantly better in quality than all other baselines across all metrics for different levels of skipping. For LLaMA-3.2-3B, again FiRST is $95\%$ in output quality of gold (base + LoRA) for 10% skipping (Table 2) with similar gains in latency (upto 16% overall) (Table 4) over the LoRA fine-tuned base model. Moreover, it is better than all other layer skipping strategies across all metrics.
Natural Questions Dataset: For the LLaMA-3-8B model, FiRST retains over $92\%$ of the EM achieved by the non-skipping gold standard (Table 2), with overall latency gains of 5-12% (Table 4). It is significantly better in quality than all other baselines for different levels of skipping. For LLaMA-3.2-3B, again FiRST is $93\%$ in output quality of gold (base + LoRA) for 10% skipping (Table 2) with gains in latency of 4-10% overall (Table 4) over the LoRA fine-tuned base model. Moreover, it is better than all other layer skipping strategies. While more sophisticated retrieval strategies could further enhance performance, our goal was to demonstrate that FiRST maintains quality performance and latency gains even in a retrieval-based QA setting.
@@ A.4 Training and Inference Setup
  • Training settings: We perform extensive experiments on two models, namely LLaMA-3-8B and LLaMA-3.2-3B from Meta, which consist of 32 and 28 layers, respectively. Training of routers and LoRA adapters is conducted on A100 80GB GPUs, with training/inference is performed in full precision to avoid performance degradation due to quantization. The training process employs our custom loss function and continues for a fixed number of epochs, terminating when the validation loss fails to improve over 4 consecutive steps. The learning rate is set between $1e^{-4}$ and $3e^{-4}$ - a cosine scheduler is used to adjust the learning rate. Gradients are accumulated after five steps, and we set the regularization loss coefficient ($\lambda=0.01$), ensuring it meaningfully contributes to the overall loss without overpowering primary losses like cross-entropy and penalization loss. After training, we verify the router weight norms (e.g., thresholds of $0.1$ or $0.05$) to ensure they remain stable—neither exploding nor vanishing—preventing overfitting or underfitting. For LoRA fine-tuning, we employ a rank of 8, a dropout rate of 0.1, and a scaling factor ($\text{lora\_alpha}=32$). Our approach involves two key penalization coefficients: $\alpha$ for the router and $\beta$ for LoRA training. The non-skip coefficient ($\alpha$) is adjusted based on dataset characteristics, sequence length, and model depth, requiring some tuning. Typically, router penalization ($\alpha$) is set 2–4× higher than LoRA’s ($\beta$) based on experimental observations to maintain a consistent skipping percentage across both training phases. A reasonable starting point is to tune $\alpha$ for a 10–15% skipping range, then gradually increase it at regular intervals to encourage higher skipping levels. In addition to the above hyperparameters, we define the maximum sequence length based on the task. For translation, it is set to 128 for router training and 256 for LoRA training. Similarly, for summarization, the sequence length is set to 500 and 700, respectively. For Question Answering, this length is set to 512. The prompts for the different tasks related to training / inference are shown in Appendix A.5.
@@ A.8 Generalizability and Reusability of Routers
We also investigate whether a router trained on one dataset can be applied directly to another dataset for the same task, without any additional training. Analysis of layer‐skipping patterns suggests that models processing different datasets of the same task exhibit similar internal behavior, implying that a router learned on one dataset could be reused elsewhere. To validate this idea, we train routers on English→German translation data and test them on English→Chinese translation, and likewise train on SQuAD for question answering before evaluating on the Natural Questions dataset. The quantitative results of these cross-dataset evaluations are presented in Table 13. In the QA experiments, the router trained and tested on NQ itself (denoted as Original in Table 13) and the router trained on SQuAD but tested on NQ (Cross-dataset) both skip a comparable fraction of layers. Despite being trained on different data, the SQuAD‐trained router achieves QA accuracy (EM score) very close to that of the NQ‐trained router, as detailed in Table 13. This indicates that the essential decision patterns learned by the router transfer well across QA datasets when using a similar level of layer skipping. A similar pattern emerges in machine translation. Whether trained on English→German or directly on English→Chinese data, routers that skip the same proportion of layers produce very similar translation quality metrics (BLEU and COMET in Table 13) on the English→Chinese task. This holds across different model sizes and skip settings. Together, these results demonstrate that routers learned on one dataset can be effectively reused on another dataset for the same task, as long as they are configured to skip a similar amount of computation. This reusability can lead to substantial savings in both training time and computing resources.